OpenStudy (anonymous):

A transverse wave is described by y(x,t) = 0.0065sin(3.57x – 27.7t + π/4) (Units are standard units.) Determine: a.) Amplitude b.) Wavelength c.) Frequency d.) Period e.) Speed of propagation f.) Direction of travel g.) Phase Constant h.) The y-position of a point at x = 1.50m and t = 0.12 s. (remember radians!)

OpenStudy (anonymous):

A=0.0065m wavelength=2*pi/k (k=wave number, here 3.57) frequency=w/(2pi) (w=angular frequency, here 27.7) Time period=1/frequency Speed of propagation or wave speed=frequency*wavelength Direction of travel=positive x axis (since - sign) Phase constant=pi/4 y=0.0065sin(3.57*1.50-27.7*0.12+pi/4)

OpenStudy (anonymous):

Thank you very much! That was so helpful. I appreciate you giving the explanations of how to get each answer and not just giving me the answer! Seriously, THANK YOU!

OpenStudy (anonymous):

A=.0065 m wavelength=v/f w=2(pi)/T T=1/f v= differentiation of main function and a= differentiation 0f v function . phase constant =(pi)/4 . all of them can help you to solve this problem .